# Statistical significance in choice modelling: computation, usage and reporting

### v2

@@ 2 Measures of uncertainty
Finally, analysts should also be mindful of a related issue, namely the risk of convergence to local optima that are (potentially very) inferior to the global optimum. This issue is more prevalent than most analysts might suspect (cf. Hess et al., 2025), and thus of course also adds some caveats to the faith analysts should place in the standard errors produced by their estimation software.
@@ 3 Confidence intervals
The standard way of computing CIs is to use asymptotic MLE properties, giving us asymptotic CIs. Formally, a $C$confidence interval for $\beta_{k}$ would be obtained as:
@@
where $\alpha=1-\frac{C}{100}$, such that $z^{\frac{\alpha}{2}}$ is the upper $\frac{\alpha}{2}$ critical value for a $\mathcal{N}(0,1)$ distribution, e.g., using $\hat{\beta}_{k}\pm 1.96\cdot\hat{\sigma}_{k}$ for a $95\%$ confidence interval. This process is explained in Figure 1, which highlights the property of asymptotic normality. The link to the asymptotic normality assumption is clear; with a standard Normal distribution, $\frac{\alpha}{2}$ of the distribution lies below $-z^{\frac{\alpha}{2}}$, while $\frac{\alpha}{2}$ of the distribution lies above $+z^{\frac{\alpha}{2}}$. The addition of $\hat{\beta}_{k}$ and multiplication of $z^{\frac{\alpha}{2}}$ by $\hat{\sigma}_{k}$ in Equation (10) simply moves us from a standard Normal distribution to the asymptotic Normal distribution for the MLE of $\beta_{k}$.
The calculation of asymptotic CIs in this manner is possible for individual parameters, or the computed value of functions of multiple parameters, using standard errors obtained via the Delta method (cf. Daly et al., 2012). Analysts can use classical standard errors, robust standard errors or bootstrap standard errors in Equation (10). The key question that arises at this stage, however, is the impact of the assumptions underlying the computation of asymptotic confidence intervals.
To be specific, if the estimated parameters were distributed exactly normal, this would imply that the log-likelihood function was quadratic in the parameters. A shift away from the optimum $\hat{\beta}_{k}$ for parameter $\beta_{k}$ by a value $\Delta_{\beta_{k}}$ would reduce the log-likelihood by $\Delta LL=\frac{{\Delta_{\beta_{k}}}^{2}}{2{\hat{\sigma}_{k}}^{2}}$. However, the ‘asymptotic’ qualifier means that the property applies exactly only for very large data sets. For finite data sets, it is an approximation that applies only in the neighbourhood of the optimum, i.e. for small $\Delta_{\beta_{k}}$. Indeed, as is shown by Armstrong et al. (2001), for example, the log-likelihood function is not exactly quadratic, so that reductions in likelihood do not follow this formula exactly at all. What happens as we move further away from the optimum is not defined by maximum likelihood theory. The parameter estimates are not distributed exactly normal, and any theoretical insights/claims based on an assumption of normality are misguided. The problem is that what defines “close” is unclear and depends on the model. There is thus no guarantee that the shape of a $C$confidence interval (especially for large $C$) is normal, that the bounds are at or close to $\hat{\beta}_{k}\pm z^{\frac{\alpha}{2}}\cdot\hat{\sigma}_{k}$ or that the bounds are symmetric around the MLE, $\hat{\beta}$.
@@ 4 Hypothesis tests
4.1 Formulation of hypotheses
@@
In statistical tests, an analyst forms two hypotheses, a null hypothesis, $\mathbf{H_{0}}$, and an alternative hypothesis $\mathbf{H_{1}}$. The test process then involves determining whether there is sufficient evidence to reject $\mathbf{H_{0}}$ in favour of $\mathbf{H_{1}}$. The specification of hypotheses differs depending on whether an analyst conducts tests for individual parameters or tests comparing different models.
4.1.1 Tests on individual parameters
When using statistical tests for individual parameters, the interest is in whether a given parameter differs from a specific value. We would then formulate a null hypothesis for parameter $\beta_{k}$ as:
$\mathbf{H_{0}}$:
$\beta_{k}=\beta_{k,H_{0}}$
The most common value for $\beta_{k,H_{0}}$ in this case is zero [^6] For some parameters, such as multipliers, scale parameters, or nesting parameters, the useful test is one where $\mathbf{H_{0}}:\beta_{k}=1$., which means that the test is used to determine whether $\beta_{k}$ has an effect in the model, i.e. setting:
$\mathbf{H_{0}}$:
$\beta_{k}=0$
What receives almost no attention in discussions in published work is the definition of the alternate hypothesis, i.e. $\mathbf{H_{1}}$. The default assumption is to set this as:
$\mathbf{H_{1}}$:
$\beta_{k}\neq 0$,
which equates to using a two-sided test. For confidence intervals, it clearly makes sense to work with a two-sided confidence interval, looking at the distribution to either side of the estimated value. However, a different reasoning should apply in hypothesis testing. Indeed, for the majority of parameters used in choice models, there is a strong a priori sign assumption. For example, a cost coefficient is a priori assumed to be negative. A model that produces a positive cost coefficient would be rejected on the basis of counter-intuitive results. This in turn means that, when such a strong sign assumption exists, this should be reflected in the formulation of $\mathbf{H_{1}}$. For example, if we a priori assume that $\beta_{k}$ should be negative, then we would use:
$\mathbf{H_{0}}$:
$\beta_{k}=0$,
$\mathbf{H_{1}}$:
$\beta_{k}<0$,
which equates to using a one-sided test.
Many papers report the outcomes of such tests without indicating if a one-sided or two-sided test was used, and the default assumptions in many software packages is two-sided [^7] Apollo (Hess and Palma, 2019) uses one-sided tests by default but gives the user the option to request two-sided tests instead.. The inappropriate use of a two-sided test has two implications. First, it means that the definition of a type I error would cover two areas, namely that where a parameter is positive, or where it is very negative. Second, in terms of actual findings, the use of a two-sided test will produce $p$-values that are double those of a one-sided test, increasing the risk of type II errors [^8] For more details on $p$-values and type I and type II errors, see Section 4.3..
A special case that warrants some attention is for a parameter that has a theoretical requirement to be contained between certain values, such as a structural parameter ($\lambda$) in a nested logit model that needs to be between 00 and $1$. A common test reported in papers is with:
$\mathbf{H_{0}}$:
$\lambda=1$,
$\mathbf{H_{1}}$:
$\lambda\neq 1$,
which would again be a two-sided test, which would mean that values that are far below the estimate, but potentially still above 00 would also be unacceptable. In such a situation, it would be safer for an analyst to instead conduct two one-sided tests, namely
$\mathbf{H_{0}}$:
$\lambda=1$,
$\mathbf{H_{1}}$:
$\lambda<1$,
and
$\mathbf{H_{0}}$:
$\lambda=0$,
$\mathbf{H_{1}}$:
$\lambda>0$.
As an alternative to conducting two one-sided tests, an analyst could also directly work out the $p$-value as being the probability of the estimate falling outside the $\left[0,1\right]$ interval.
4.1.2 Tests comparing models
The null hypothesis is then:
$\mathbf{H_{0}}$:
$LL_{M_{g}}=LL_{M_{r}}$.
$\mathbf{H_{1}}$:
$LL_{M_{g}}>LL_{M_{r}}$,
meaning that a one-sided test will be used.
In case one needs to test non-nested hypothesis, cases in which one model is not a restricted version of the other, the reader is referred to the work of Ben-Akiva and Swait (1986), who determined a bound for the probability that a model may have a larger adjusted $\bar{\rho}^{2}$, while not being the one behind the data generation process [^9] We define $\bar{\rho}^{2}=1-\frac{LL_{\hat{\beta}}-K}{LL_{0}}$, where $LL_{\hat{\beta}}$ and $LL_{0}$ are the log-likelihoods at convergence and for a model with equal probabilities for all available alternatives, respectively, while $K$ is the number of estimated parameters (cf. Domencich and McFadden, 1975; Mokhtarian, 2016)..
4.2 A trinity of tests
$\mathbf{H_{0}}$:
$\beta_{k}=0$,
$\mathbf{H_{1}}$:
$\beta_{k}\neq 0$.
Likelihood ratio (LR) test:
The LR test compares the model fit of a general model and its constrained version. If $H_{0}$ is true, the test statistic converges asymptotically to a $\chi^{2}$ distribution with $d$ degrees of freedom, where $d$ is the number of additional parameters in the general model. Let $\hat{\beta}$ be the parameters obtained in the unconstrained estimation, while $\tilde{\beta}$ refers the constrained case, e.g. where we constrain $\beta_{k}=0$. We then have:
Wald test:
where, with the above, $\beta_{k,H_{0}}=0$.
Given the discussions in Daly et al. (2012) in relation to the MLE properties of transformations of parameters, the same test can also be used for functions of parameters with standard errors obtained via the Delta method.
Lagrange multiplier (LM) test1212footnotemark: 12:
To understand the relationship between the three tests described above, Figure 2 is helpful. Of the three tests, only the LR test uses information from both the general and constrained model by comparing $LL\left(\hat{\beta}\right)$ and $LL\left(\tilde{\beta}\right)$. The Wald test, and thus also the widely used $t$-ratios, uses information only from the general model, in terms of $\hat{\beta}$ and $\hat{\sigma}$. Finally, the LM test uses information only on the shape of the unconstrained log-likelihood function at the solution from the constrained model $\tilde{\beta}$, indicated by the red slope line. For the LR test, both models thus need to be estimated, for the Wald test, only the general model, and for the LM test, only the constrained model. The three tests are asymptotically equivalent but may diverge substantially with smaller samples. Overall, by using both log-likelihood values, the LR test does not make assumptions about the shape of the likelihood function, i.e. it does not make any asymptotic assumptions. If both the restricted and the unrestricted models are easy to evaluate, it should be preferred as it generally has more power (see e.g. Gudicha et al., 2016). This test is often preferred in DCM when assessing constraints that involve many coefficients together, since the likelihood of the restricted and the unrestricted model can be directly obtained from the model output, but in the context of behaviourally important individual parameters, it should potentially also be the preferred test when looking at individual parameters.
4.3 Making decisions based on tests
In Section 4.1, we discussed the process of formulating a null and alternate hypothesis, while in Section 4.2, we described the three key tests used to decide whether the null hypothesis $\mathbf{H_{0}}$ can be rejected given the evidence found for $\mathbf{H_{1}}$.
@@ 4.3 Making decisions based on tests
As shown in Table 2, there are four possible cases that combine the truth in the data with the decision by the analyst. The analyst does not know whether $\mathbf{H_{0}}$ is true or false, and thus needs to determine what risk of making a wrong decision is acceptable. The core emphasis in statistical tests conducted in choice modelling relates to how confident we can be in rejecting $\mathbf{H_{0}}$, and in particular what the risk of a type I error is. When conducting a statistical test, we obtain a $p$ value, which is the probability of observing a value as high as the test statistic when $\mathbf{H_{0}}$ is true. In simpler terms, the $p$-value is the probability that our finding in support of rejecting $\mathbf{H_{0}}$ was obtained by chance. A smaller $p$-value suggests stronger evidence to reject $\mathbf{H_{0}}$. An analyst needs to set the maximum acceptable probability of making a Type I error, where this is defined as $\alpha_{\text{type I}}$, also known as the significance level of the test [^14] We add the subscript in $\alpha_{\text{type I}}$ to avoid any confusion with model parameters defined as $\alpha$.. The $p$-value obtained via a statistical test is compared to $\alpha_{typeI}$, and $\mathbf{H_{0}}$ is rejected if $p\leq\alpha_{typeI}$. Setting a lower $\alpha_{typeI}$ thus reduces the risk of a type I error.
@@
The point about $\alpha_{\text{type I}}$ being the significance of a test also highlights a common bad practice. Analysts routinely describe the findings of statistical tests as showing that a parameter is “significantly different from 00”. It should first be noted that hypothesis testing is only concerned with whether the null hypothesis can be rejected, not whether the alternative hypothesis is true. In other words, it establishes whether our finding could be due to chance and that the true value is in fact zero (if $\mathbf{H_{0}}$: $\beta_{k}=0$). What is commonly described as a test of significance is thus a test to see whether the null hypothesis of a parameter being equal to zero can be rejected. The statistical significance of the test relates to the probability of $H_{0}$ being rejected when it is in fact true, i.e. a type I error. It is thus incorrect to talk about a parameter being $95\%$ significant. The significance level in this case is $5\%$, and we can reject $H_{0}$ with a $95\%$ level of confidence, not significance. The relationship with the earlier discussion on confidence intervals is clear. Using a negative estimate of $\hat{\beta}$ as our example, if we can reject $H_{0}:\beta=0$ with a confidence level of $C\%$ for a two-sided test, then this means that 00 forms the upper bound of the $C\%$ confidence interval around the estimated value $\hat{\beta_{k}}$.
It is also important to note that the definition of $\alpha_{\text{type I}}$ depends on the parametric sampling distribution, which varies by the type of test, holds asymptotically for DCM, and relies on specific assumptions. Consequently, the true (empirical) $\alpha_{\text{type I}}$ may differ from the nominal $\alpha_{\text{type I}}$. This difference is referred to as the size distortion of the statistical test. If the distortion is positive or “liberal”, the empirical $\alpha_{\text{type I}}$ is larger than the nominal one, and this increases the likelihood of committing more Type I errors than expected. Conversely, if the distortion is negative or “conservative”, the test makes fewer Type I errors than planned, but becomes more prone to Type II errors. Size distortion can be analyzed using Monte Carlo simulations, which replicate the resampling process and provide a useful tool for comparing the performance of different statistical tests.

### v3

@@ 2.6 Identification and over-specification issues
Finally, analysts should also be mindful of a related issue, namely the risk of convergence to local optima that are (potentially very) inferior to the global optimum. This issue is more prevalent than most analysts might suspect (cf. Hess et al., 2025), and thus of course also adds some caveats to the faith analysts should place in the standard errors produced by their estimation software. Specifically, inference may reflect curvature at a local optimum rather than global likelihood geometry. Given the resulting possibility of substantially different estimates and standard errors, this could mean not only different effect sizes but a likelihood that with one local optimum, we find an effect, while with another, we do not, leading to different conclusions. A detailed discussion of this topic is beyond the scope of the present paper, but analysts are encouraged to invest time in investigating the use of different starting values, robust estimation approaches and targeted solutions (see e.g. Bierlaire et al., 2010).
@@ 3 Confidence intervals
The standard way of computing CIs is to use asymptotic MLE properties, giving us asymptotic CIs. Formally, a CC% confidence interval for $\beta_{k}$ would be obtained as:
@@
where $\alpha=1-\frac{C}{100}$, such that $z^{\frac{\alpha}{2}}$ is the upper $\frac{\alpha}{2}$ critical value for a $\mathcal{N}(0,1)$ distribution, e.g., using $\hat{\beta}_{k}\pm 1.96\cdot\hat{\sigma}_{k}$ for a $95\%$ confidence interval. This process is explained in Figure 1, which highlights the property of asymptotic normality. The link to the asymptotic normality assumption is clear; with a standard Normal distribution, $\frac{\alpha}{2}$ of the distribution lies below $z^{\frac{\alpha}{2}}$, while $\frac{\alpha}{2}$ of the distribution lies above $z^{\frac{\alpha}{2}}$. The addition of $\hat{\beta}_{k}$ and multiplication of $z^{\frac{\alpha}{2}}$ by $\hat{\sigma}_{k}$ in Equation (11) simply moves us from a standard Normal distribution to the asymptotic Normal distribution for the MLE of $\beta_{k}$.
The calculation of asymptotic CIs in this manner is possible for individual parameters, or the computed value of functions of multiple parameters, using standard errors obtained via the Delta method (cf. Daly et al., 2012). If analysts wish to consider confidence intervals around the actual distribution of a random coefficient (or ratio of coefficients), then they further need to consider the work of Bliemer and Rose (2013); Scaccia et al. (2023).
In computing confidence intervals, analysts can use classical standard errors, robust standard errors or bootstrap standard errors in Equation (11). The key question that arises at this stage, however, is the impact of the assumptions underlying the computation of asymptotic confidence intervals.
To be specific, if the estimated parameters were distributed exactly normal, this would imply that the log-likelihood function was quadratic in the parameters. A shift away from the optimum $\hat{\beta}_{k}$ for parameter $\beta_{k}$ by a value $\Delta_{\beta_{k}}$ would reduce the log-likelihood by $\Delta LL=\frac{{\Delta_{\beta_{k}}}^{2}}{2{\hat{\sigma}_{k}}^{2}}$. However, the ‘asymptotic’ qualifier means that the property applies exactly only as the sample size tends towards infinity. For finite data sets, it is an approximation that applies only in the neighbourhood of the optimum, i.e. for small $\Delta_{\beta_{k}}$. Indeed, as is shown by Armstrong et al. (2001), for example, the log-likelihood function is not exactly quadratic, so that reductions in likelihood do not follow this formula exactly at all. What happens as we move further away from the optimum is not defined by maximum likelihood theory. The parameter estimates are not distributed exactly normal, and any theoretical insights/claims based on an assumption of normality are misguided. The problem is that what defines “close” is unclear and depends on the model. There is thus no guarantee that the shape of a CC% confidence interval (especially for large $C$) is normal, that the bounds are at or close to $\hat{\beta}_{k}\pm z^{\frac{\alpha}{2}}\cdot\hat{\sigma}_{k}$ or that the bounds are symmetric around the MLE, $\hat{\beta}$.
@@ 4 Hypothesis tests
4.1 Formulation of hypotheses and possible test outcomes
@@
In statistical tests, an analyst forms two hypotheses, a null hypothesis, $\mathbf{H_{0}}$, and an alternative hypothesis $\mathbf{H_{1}}$. The test process then involves determining whether there is sufficient evidence to reject $\mathbf{H_{0}}$ in favour of $\mathbf{H_{1}}$.
@@
As shown in Table 2, there are four possible cases that combine the truth in the data with the decision by the analyst. The analyst does not know whether $\mathbf{H_{0}}$ is true or false, and thus needs to determine what risk of making a wrong decision is acceptable. The core emphasis in statistical tests conducted in choice modelling relates to how confident we can be in rejecting $\mathbf{H_{0}}$, and in particular what the risk of a type I error is, i.e. rejecting the null hypothesis even though it is true. When conducting a statistical test, we obtain a $p$ value for the given test statistic, which is the probability of observing a value as high as the test statistic when $\mathbf{H_{0}}$ is true. In simpler terms, the $p$-value is the probability that our finding in support of rejecting $\mathbf{H_{0}}$ was obtained by chance. A smaller $p$-value suggests stronger evidence to reject $\mathbf{H_{0}}$. An analyst needs to set the maximum acceptable probability of making a Type I error, where this is defined as $\alpha_{\text{type I}}$, also known as the significance level of the test [^5] We add the subscript in $\alpha_{\text{type I}}$ to avoid any confusion with model parameters defined as $\alpha$.. The $p$-value obtained via a statistical test is compared to $\alpha_{typeI}$, and $\mathbf{H_{0}}$ is rejected if $p\leq\alpha_{typeI}$. Setting a lower $\alpha_{typeI}$ thus reduces the risk of a type I error.
@@
As described above, a statistical test is used to determine whether the data support $\mathbf{H_{1}}$ over $\mathbf{H_{0}}$, i.e. allow us to reject $\mathbf{H_{0}}$. There are some important nuances that are often misunderstood. First, rejecting $\mathbf{H_{0}}$ does not mean accepting $\mathbf{H_{1}}$. It simply means that $\mathbf{H_{0}}$ is unlikely to be true, given the evidence in support of $\mathbf{H_{1}}$. Second, a failure to reject $\mathbf{H_{0}}$ does not mean that $\mathbf{H_{0}}$ is true, but simply means that we do not have strong enough evidence to reject $\mathbf{H_{0}}$. Formally, this means that analysts should not make statements that they have evidence that an effect exists, or even that they are testing whether an effect exists. Indeed, the correct statement would be that they can reject the hypothesis that an effect does not exist. Another related point concerns was Ziliak and McCloskey (2008) call the fallacy of the transposed conditional. A statistical test considers the probability of observing the data given that $\mathbf{H_{0}}$ is true, expressed as the $p$-value. This is not the same as the probability that $\mathbf{H_{0}}$ is true given the data.
The specification of hypotheses differs depending on whether an analyst conducts tests for individual parameters or tests comparing different models, as we will discuss in the following sections.
4.2 Hypotheses for tests on individual parameters
When using statistical tests for individual parameters, our key interest is in whether a given parameter differs from a specific value. We first formulate a null hypothesis for parameter $\beta_{k}$ as:
$\mathbf{H_{0}}$:
$\beta_{k}=\beta_{k,H_{0}}$
Notwithstanding the discussions in the introduction about the widespread question of how valid it is to consider the possibility of a parameter having a value of zero, the most common value for $\beta_{k,H_{0}}$ in this case is indeed zero [^7] For some parameters, such as multipliers, scale parameters, or nesting parameters, the useful test is one where $\mathbf{H_{0}}:\beta_{k}=1$.. There are of course parameters where it is entirely reasonable to consider the possibility that the estimate is zero. For many others, such as key behavioural attributes like time and cost, this is more doubtful. Even here though, reasons exist, for example the possibility that in the case of poor data or inappropriate design for a stated choice survey, the attribute has little or no possibility of affecting behaviour. In this case, an analyst will want to be able to test this, which means that the test is used to determine whether $\beta_{k}$ has an effect in the model. It is important to recognise that, formally, we do not establish whether an effect exists, but whether we can reject the null hypothesis that no effect exists. Specifically, we thus set:
$\mathbf{H_{0}}$:
$\beta_{k}=0$
Alongside the null hypothesis, there is a requirement for the definition of the alternate hypothesis, i.e. $\mathbf{H_{1}}$. With generally limited discussion in published work, the default assumption is to set this as:
$\mathbf{H_{1}}$:
$\beta_{k}\neq 0$,
which equates to using a two-sided test. For confidence intervals, it clearly makes sense to work with a two-sided confidence interval, looking at the distribution to either side of the estimated value. However, a different reasoning should apply in hypothesis testing. Indeed, for the majority of parameters used in choice models, there is a strong a priori sign assumption. For example, a cost coefficient is a priori assumed to be non-positive. A model that produces a positive cost coefficient would be rejected on the basis of counter-intuitive results. This in turn means that, when such a strong sign assumption exists, this should be reflected in the formulation of $\mathbf{H_{1}}$. For example, if we a priori assume that $\beta_{k}$ should be non-positive, then we would use:
$\mathbf{H_{0}}$:
$\beta_{k}=0$,
$\mathbf{H_{1}}$:
$\beta_{k}<0$,
which equates to using a one-sided test.
Many papers report the outcomes of such tests without indicating if a one-sided or two-sided test was used, and the default assumptions in many software packages is two-sided [^8] Apollo (Hess and Palma, 2019) uses one-sided tests by default but gives the user the option to request two-sided tests instead.. The inappropriate use of a two-sided test has two implications. First, it means that the definition of a type I error would cover two areas, namely that where a parameter is positive, or where it is very negative. Second, in terms of actual findings, the use of a two-sided test will produce $p$-values that are double those of a one-sided test, increasing the risk of type II errors.
A special case that warrants some attention is for a parameter that has a theoretical requirement to be contained between certain values, such as a structural parameter ($\lambda$) in a nested logit model that needs to be between 0 and $1$. A common test reported in papers is with:
$\mathbf{H_{0}}$:
$\lambda=1$,
$\mathbf{H_{1}}$:
$\lambda\neq 1$,
which would again be a two-sided test, which would mean that values that are far below the estimate, but potentially still above 0 would also be unacceptable. In such a situation, it would be safer for an analyst to instead conduct two one-sided tests, namely
$\mathbf{H_{0}}$:
$\lambda=1$,
$\mathbf{H_{1}}$:
$\lambda<1$,
and
$\mathbf{H_{0}}$:
$\lambda=0$,
$\mathbf{H_{1}}$:
$\lambda>0$.
As an alternative to conducting two one-sided tests, an analyst could also directly work out the $p$-value as being the probability of the estimate falling outside the $\left[0,1\right]$ interval, i.e. what part of a $N\left(\hat{\lambda},\hat{\sigma}_{\lambda}\right)$ falls outside this interval, where $\hat{\lambda}$ and $\hat{\sigma}_{\lambda}$ are the MLE and standard error, respectively.
4.3 Hypotheses for tests comparing models
The null hypothesis is then:
$\mathbf{H_{0}}$:
$LL_{M_{g}}=LL_{M_{r}}$.
$\mathbf{H_{1}}$:
$LL_{M_{g}}>LL_{M_{r}}$,
meaning that a one-sided test will be used.
4.4 A trinity of tests
$\mathbf{H_{0}}$:
$\beta_{k}=0$,
$\mathbf{H_{1}}$:
$\beta_{k}\neq 0$.
The three tests of interest are:
Likelihood ratio (LR) test:
The LR test compares the fit of a general model and its constrained version. If $H_{0}$ is true, the test statistic converges asymptotically to a $\chi^{2}$ distribution with $d$ degrees of freedom, where $d$ is the number of additional parameters in the general model. Let $\hat{\beta}$ be the parameters obtained in the unconstrained estimation, while $\tilde{\beta}$ refers the constrained case, e.g. where we constrain $\beta_{k}=0$. We then have:
Wald test:
where, with the above, $\beta_{k,H_{0}}=0$.
Lagrange multiplier (LM) test1111footnotemark: 11:
[…]
4.5 Considerations specific to choice models
Our discussion so far has looked at the topic from a broad perspective, and we now turn our attention to a few considerations that are specific to choice models. We again look separately at tests for individual measures, and at tests relating to model comparisons.
4.5.1 Tests for derived measures
The discussion in Section 4.2 and the details on the Wald test (and $t$-ratios in Section 4.4) focussed entirely on individual parameters. We now extend this to functions of parameters, and finally the incorporation of random heterogeneity.
The first extension is straightforward, given the properties of the Delta method, as discussed by Daly et al. (2012), and as highlighted in Section 2.4. An analyst can thus use the exact same formulation of hypotheses and tests discussed here for the case of measures computed on the basis of estimated model parameters, such as for example a MRS computation.
The second extension, i.e. the presence of random heterogeneity, introduces additional complexity. Let us again focus on the case of a continuous mixture model, such that $\beta_{1}$ follows a distribution across individual people. The formulation of hypotheses in this case can take multiple forms. An analyst can conduct tests on the parameters representing the distribution, such as for example the mean and standard deviation. A test on the former would look at whether we can reject the hypothesis that the mean sensitivity is zero, while a test on the latter would allow us to test whether we can reject the hypothesis that there is no heterogeneity across individuals. This is however different from formulating a test on a distributed $\beta_{1}$, i.e. testing whether the population level distribution is different from zero. For this, an analyst needs to consider both the parameter uncertainty and the distribution of $\beta_{1}$ across people, for which we again refer the reader to Bliemer and Rose (2013); Scaccia et al. (2023).
4.5.2 Model specification tests
Choice modellers routinely estimate multiple different specifications on the same data and compare the resulting models against each other. This relates to situations where additional terms are added to a model, for example interactions between variables, treatments of non-linearity, heterogeneity, and more flexible assumptions about the error structure, notably in relation to correlation between the unobserved utility components for different alternatives.
To a large extent, such comparisons can make use of LR tests as they compare a restricted version of a model to a more general model. This includes for example models with additional interaction terms, the comparison of models with more flexible error structures to simpler models (e.g. Nested vs Multinomial Logit), or models with and without random heterogeneity (comparing Mixed Logit to Multinomial Logit). In each case, a simple parameter constraint would make the general model equivalent to the restricted model (generally setting concerned parameters to 0, or in the case of nesting structures and multipliers to $1$)
At other times, a LR test is not possible, i.e. when one model is not a restricted version of the other model that can be explained by a simple parameter constraints. Some of these examples are well known, such as when comparing models with different types of non-linearity (e.g. linear vs logarithmic), models with different treatments of heterogeneity (e.g. Mixed Logit vs Latent Class) or models from different model families (e.g. Random Utility vs Decision Field Theory). There are also examples that are less well recognised, and where analysts may mistakenly think they can apply a LR test. One example relates to comparing Latent Class models with different numbers of classes. A three class model can collapse to a two class model either by one class getting a zero class allocation probability, or by the parameters inside two classes collapsing to the same values. Another example relates to comparing a Cross-Nested Logit model to a Nested Logit model. Again, multiple restrictions potentially yield the simpler model, namely restrictions on the nesting parameters or the allocation parameters In both those cases, there are multiple ways in which a model can collapse into the other and hence the LR-test does not make sense, as it is not clear what the appropriate degrees of freedom would be.
These situations relate to testing non-nested hypotheses. Informal comparisons can be made on the basis of goodness of fit statistics such as the adjusted $\bar{\rho}^{2}$ measure (cf. Domencich and McFadden, 1975; Mokhtarian, 2016), or the Akaike Information Criterion (AIC, cf. Akaike, 1974) and the Bayesian Information Criterion (BIC, cf. Schwarz, 1978).
The adjusted $\bar{\rho}^{2}$ has special interest given that a formal test exists using it for the comparison of non-nested models. Let us define $\bar{\rho}^{2}=1-\frac{LL_{\hat{\beta}}-K}{LL_{0}}$, where $LL_{\hat{\beta}}$ and $LL_{0}$ are the log-likelihoods at convergence and for a model with equal probabilities for all available alternatives, respectively, while $K$ is the number of estimated parameters. Ben-Akiva and Swait (1986) determined a bound for the probability that a model may have a larger adjusted $\bar{\rho}^{2}$, while not being the one behind the data generation process, given by:
$\texttt{Pr}(\bar{\rho}_{1}^{2}-\bar{\rho}_{2}^{2}\geq z)\leq\Phi\left[-\left(2\cdot z\cdot LL\left(\beta_{ES}\right)+df_{1}-df_{2}\right)^{\frac{1}{2}}\right],$ (16)
where $z$ is observed difference in adjusted $\rho^{2}$, $\Phi$ is cumulative standard normal distribution, $LL\left(\beta_{ES}\right)$ is log-likelihood at equal shares, and $df_{1}$ and $df_{2}$ are the number of parameters for models 1 and 2.
This repertoire of LR tests, goodness of fit statistics, and the Ben-Akiva & Swait test cover the majority of cases of model comparisons in choice modelling. The well known Swait & Louviere test (Swait and Louviere, 1993) for example uses a sequence of separate LR tests for comparing a generic pooled model to a model with scale differences and separate models. Similarly, comparing models that make different error structure assumptions to test for violations of the IIA property [^13] Note this is different from comparing parameter estimates of models estimated with a full set of alternatives as opposed to a subset, as in the Hausman and McFadden (1984) work, albeit that there also, statistical tests could then be used to compare estimates, at least in the form of MRS. relies on LR tests (cf. Brownstone and Train, 1999).
4.6 Making decisions based on tests